Having trouble with #5Saratoga Company manufactures jobs to customer specifications. The company is conducting a time-driven activity-based costing
Answer #1

Calculation of Employee costs-

Particulars Amounts
No.of employees 9
Average salary per employee $ 28000
Total salary for the year $ 252000
Weeks of the employment per year 52
Hours per week / employee 40
Total No. of hours/ employee p.a 2080
Total No. of hours of all the employees for the year 18720
Practical capacity % 85%
Usable capacity 15973.50
Cost / minute (252000/(18720*60)) 0.23

2) Calculating cost for activities required:

Particulars Job X Job Y Job Z Total
Requisitions Processed 10 7 6
Time for requisition processing @ 15 mins. 150 105 90 345
Bid Evaluations 2 1 3
Time Required for Bid Evaluation @ 45 mins 90 45 135 270
No. of Inspections 5 1 5
Time Required for Inspection @ 30 mins 150 30 150 330
Total Time Required (Mins.) 390 180 375

3) Based on revised activity demands - values are calculated as follows:

Particulars Requisition Processing Bid Evaluation Inspection Total
Total Demand (a) 7100 8650 10500
Mins/Unit (b) 15 45 30
Total Time (Mins.) = (c=a*b) 106500 389250 315000 810750
Practical capacity (d) 85% 85% 85%

Total Time required

(e=c/d)

125294.12 457941.18 370588.24
Total Hours Required (f=e/60) 2088.23 7632.35 6176.47 15897.15
Total No. of employees required = Hours demanded / 2080 (g) 7.64 i.e. 8 emloyees.
Total cost of 8 employees (h =g*28000) 224000
Allocation of cost between activities (i=h/15897.15*f) 29424.64 107545.22 87030.14 224000
Cost/Hour (j=i/f) $ 14.09 $14.09 $14.09
Cost per minute (k = j/60) $ 0.23 $ 0.23 $ 0.23

3) Impact on expenses of matching capacity with demand:

Earlier total cost were = $ 252000

Now revised costs are = $ 224000

Hence, cost has reduced as a result of matching of capacity with demands.

4) Considering the revised activity demands labour costs assigned to each of the three activities:

Particulars Job X Job Y Job Z Total
Requisitions Processed 10 7 6
Time for requisition processing @ 15 mins. 150 105 90 345
Bid Evaluations 2 1 3
Time Required for Bid Evaluation @ 45 mins 90 45 135 270
No. of Inspections 5 1 5
Time Required for Inspection @ 30 mins 150 30 150 330
Total Time Required (Mins.) 390 180 375
Cost / Min $0.23 $0.23 $0.23
Total cost $89.70 $41.40 $86.25

2) Using the revised activity demands will not affect labour costs assigned to Job X, Y and Z since using the revised demand has not either reduced the no of activities used per unit of product X,Y and Z neither has the labour rate per hour has changed.

3) Using the revised capacity Used capacity in minutes is 810750 mins i.e. 13512.50 hrs. (calculated in table above)

4) Using the revised capacity unused capacity is = (8*52*40 - 13512.50 = 2487.50 Hours).